OpenStudy (anonymous):

find the standard form of the hyperbola with v( plus minus 1, 0) through point (2, 13)

OpenStudy (amistre64):

vertices are (-1,0), and (1,0); its a convenient thing this is along the x axis i spose

OpenStudy (amistre64):

(2)^2 - (13)^2 = 1 is our goal i think

OpenStudy (amistre64):

4 196 --- - --- = 1 hmmm a^2 b^2

OpenStudy (amistre64):

a^2 = 1 since 1/1 = 1 for out vertice

OpenStudy (amistre64):

4 - 196/b^2 = 1 4b^2 -196 = b^2 3b^2 -196 = 0 b^2 = 196/3 x^2 - 3y^2/196 = 1 is te best i can do

OpenStudy (amistre64):

lol.... 13^2 = 169 ..... i always gets those confused

OpenStudy (amistre64):

\[x^2 - \frac{3y^2}{169}=1\] should be better